OpenStudy (anonymous):

find the sum of the infinite geometric series -8+4-2+1-

OpenStudy (anonymous):

\[-8+4-2+1-\cdots~~?\]

OpenStudy (anonymous):

Each successive term has an alternating sign and is half of the previous term, making the common ratio \(-\dfrac{1}{2}\). The sum of an infinite geometric series \(\displaystyle\sum_{k=0}^\infty a\times r^k=\frac{a}{1-r}\) for \(|r|<1\).

OpenStudy (anonymous):

common factor r is -1/2 so by infinite GM formula of a/(1-r)=-8/(1-(-1/2))=-8/(3/2)=-16/3

OpenStudy (anonymous):

so how would you show work for this and show your steps to finding the answer?

OpenStudy (cwrw238):

the first term = a = -8 and common ratio = r = -0.5 sum to infinity = a / (1 - r) so plug a = 8 and r -= -0.5 into the above formula